OpenStudy (minecraft40952):

Find two numbers with a sum of –46 and a difference of 4. (1 point) 25 and 21 24 and 20 –21 and –25 –22 and –26

OpenStudy (minecraft40952):

i will fan anyone who can Give the answer and explain how you got it.

OpenStudy (minecraft40952):

and a medal

OpenStudy (clongoria2):

Just add the options together 25 + 21 = 46 And so on... Can you give the answer?

OpenStudy (misty1212):

HI Check and see which pair works !

OpenStudy (minecraft40952):

i dont understand

OpenStudy (misty1212):

only one pair adds so \(-46\)

OpenStudy (minecraft40952):

theres only 1 option

OpenStudy (clongoria2):

Ok so you can elimnate A and B because they are positive numbers. So C and D are left

OpenStudy (minecraft40952):

only 1 pair

OpenStudy (clongoria2):

-21 + -25 =? -22 + -26 =?

OpenStudy (clongoria2):

which one equals -46?

OpenStudy (minecraft40952):

-22+-26

OpenStudy (clongoria2):

-22 + -26 = -48 When adding TWO negative numbers just ignore the negative sign After you add them just put the negative sign back

OpenStudy (clongoria2):

So -21 + -25 | 21 + 25 = 46 = -46

OpenStudy (clongoria2):

C is your correct answer

OpenStudy (minecraft40952):

no d

OpenStudy (minecraft40952):

do -22+-26

OpenStudy (minecraft40952):

you get -46

OpenStudy (clongoria2):

Are you sure you are calculating that right? 22 + 26 _____ 48

OpenStudy (minecraft40952):

its -22 and -26

OpenStudy (clongoria2):

Calculate 21 + 25. What do you get?

OpenStudy (minecraft40952):

and im using a calculator

OpenStudy (minecraft40952):

You need it negitive

OpenStudy (minecraft40952):

read the question

OpenStudy (clongoria2):

It is C

OpenStudy (clongoria2):

I know. -21 + -25 = -46

OpenStudy (minecraft40952):

No it is not It says you need NEGATIVE 46

OpenStudy (clongoria2):

I know -21 + -25 = NEGATIVE46

OpenStudy (clongoria2):

the - sign is negative -21 + -25 = -46 -46 <---- NEGATIVE

OpenStudy (minecraft40952):

oh sorry my calculator was on a differnt setting

OpenStudy (clongoria2):

.-.

OpenStudy (minecraft40952):

Lol

OpenStudy (minecraft40952):

well there im a fan and you got a medal

OpenStudy (clongoria2):

Good luck on the rest of the test....